Match a load impedance of 100 ‐ j100 Ω to a 50 Ω transmission line using a parallel inductor (next to
the  load)  and  a  series  capacitor.  Calculate  component  values  at  1  GHz  using  Smith  Chart.  Show  all
calculations on Smith Chart.
Answer #1

mpedance with 11.65nH parallel inductor 1.8pE series capacitorGlven 2 = 10092 - 1-1002 Load impedance z = 5092 Characterstic impedance Let inductor which needs to be added in parallel hasx= 73.222 now x is inductor impedance at 1GHz 2-1-F-L = x solving for L L=245 = 27.7042 = 11.65mH Adding this inductor ZL_new1 match -= 1.84pF 2-7-1GHz-86.6